College Athletic Department sells both football and basketball tickets to

student fans. Sports fans among the Friends University student body can be classified

into three “types,” with the following reservation prices for tickets:

Reservation Prices ($)

Buyer Type

Number of Students

Football Tickets

Basketball Tickets

Football Fans

10,000

100

40

“All Sports” Fans

2,000

75

75

Basketball Fans

3,000

50

105

If the tickets for football and basketball games are sold separately, find the Athletic Department’s profit maximizing price to charge.

a. Pmax for Football tickets (Price)

b. Pmax for Basketball tickets (Price)

c. What is the maximum profit that can be achieved pricing separately?

If football and basketball tickets are sold as a bundle, find the Athletic Department’s profit maximizing price and the resulting profit.

d. Pmax(Bundle) (Price of the bundle)

e. What is the maximum profit that can be achieved pricing as a bundle?

f. Which of the two strategies should the athletic department choose? (Bundling or Separate)
Answer #1

Solution :

(a) :- Pmax for Football tickets = $100

This is because if the AD charges $75 to include all fans ; the revenue will be (75×12,000)= $900,000 which is less than the revenue. In case tickets are charged $100.

(b) :- Similarly, Pmax for Basketball tickets = $105

This is also revenue maximizing price for AD in case of basket ball.

(c) :- Maximum profit = (100×10,000)+(105×3000) = $1,315,000

(d) :- Price of bundle will be such that the AD can include all types of fan then it is a price in all groups will play.

So, Price of bundle = $75

(e) :- Maximum profit with bundle = 75×15000 = $1,125,000

(f) :- As max profit without bundle is greater than maximum profit with bundle. The AD must choose Seprate Pricing Strategy.